OpenStudy (owlcoffee):

Need some help with this limit:

OpenStudy (owlcoffee):

\[\lim_{x \rightarrow \infty} \frac{ \ln (x ^{2}+x)-\ln (x ^{2}-3x+2) }{ e ^{\frac{ x+1 }{ x ^{2} }}-e ^{\frac{ 1 }{ x }} }\]

OpenStudy (bahrom7893):

I'm thinking l'hopital's is prolly gonna help.. But we can prolly simplify this too..

OpenStudy (owlcoffee):

I was told to not use L'hopital :(

OpenStudy (bahrom7893):

simplification it is!... ln(A/B)=lnA-lnB

OpenStudy (anonymous):

It appears to approach 0.

OpenStudy (bahrom7893):

Uhm let's see.. Idk how far I'll get with this: let's just focus on the top \[ln(x^2+x)-ln(x^2-3x+2)=ln(\frac{x(x+1)}{(x-1)(x-2)})\]

OpenStudy (anonymous):

The denominator is just e^(1/x^2)) which approaches 1 as x approaches infinity.

OpenStudy (anonymous):

The numerator is what bahrom7893 has...and that clearly approaches ln(1) = 0...so you have 0/1 = 0

OpenStudy (bahrom7893):

Actually the bottom approaches 0, no?

OpenStudy (anonymous):

no. 1/x^2 approaches 0, but e^0 approaches 1.

OpenStudy (bahrom7893):

\[e^{(x+1)/x^2}\] that is \[e^0\] as x goes to infinity

OpenStudy (bahrom7893):

e^0 - e^0 = 0

OpenStudy (owlcoffee):

\[\lim_{x \rightarrow \infty}\frac{ \ln (\frac{ x ^{2}+x }{ x ^{2}-3x+2 }) }{ e ^{\frac{ 1 }{ x }}(e ^{\frac{ x+1 }{ x ^{2} }-\frac{ 1 }{ x }}-1) }\]

OpenStudy (anonymous):

Exacuse me, but the denominator is equivalent to just e^(1/x^2)) which approaches 1 as x approaches infinity.

OpenStudy (bahrom7893):

I don't see why though?

OpenStudy (anonymous):

Look at the original problem...that expression in the denominator is equivalent to e^(1/x^2)) by using the law of exponents, e^x/e^y = e^(x-y) ..use it backwards.

OpenStudy (bahrom7893):

Dude it's a minus

OpenStudy (anonymous):

(x+1)/x^2 - 1/x

OpenStudy (bahrom7893):

You can't use the law of exponents there..

OpenStudy (owlcoffee):

I think it would be a good idea to take as common factor e^1/x

OpenStudy (anonymous):

Yes, you can..by rewriting it as one exponent!

OpenStudy (bahrom7893):

\[\huge{e ^{\frac{ x+1 }{ x ^{2} }}-e ^{\frac{ 1 }{ x }}}\]IS NOT \[\huge{e ^{\frac{ x+1 }{ x ^{2} }-\frac{ 1 }{ x }}}\]

OpenStudy (owlcoffee):

oh, no no, I agree but I took it as common factor, not substracting it.

OpenStudy (bahrom7893):

Nah i'm just pointing out why the bottom approaches 0, not 1

OpenStudy (bahrom7893):

hmm taking ln is useless and so is raising to the e..... gah we need to figure out what top approaches..

OpenStudy (bahrom7893):

top is gonna approach 0

OpenStudy (bahrom7893):

the fraction inside ln is going to 1..

OpenStudy (anonymous):

Sorry...my error for the denominator.

OpenStudy (bahrom7893):

oh man this is so ideal for l'hopital..... darn it. Np, i thought i was missing something too.

OpenStudy (owlcoffee):

You factored the polynomials inside the Ln?

OpenStudy (anonymous):

Can't the first expression in the denominator be facored out by e^(1/x)?

OpenStudy (bahrom7893):

yea i did but it didn't really help, and neither did factoring out e^(1/x).. this isn't even calc at this point, it's just algebra..

OpenStudy (bahrom7893):

Actually factor it out, yea

OpenStudy (anonymous):

Isnt it e^(1/x) { e^(1/x^2) + 1}

OpenStudy (bahrom7893):

no it's not. there's a minus in the original

OpenStudy (bahrom7893):

lol too tired to think right now..

OpenStudy (anonymous):

So e^(1/x) was factored out with a - instead of our +

OpenStudy (owlcoffee):

I have an idea, I'll work on the denominator: \[e ^{\frac{ 1 }{ x }}(e ^{\frac{ x+1 }{ x ^{2} }-\frac{ 1 }{ x }}-1)\] I know that there's a equivalent: \[e ^{F(x)}-1= f(x)\] \[f(x)\rightarrow0\] so it would be: \[e ^{\frac{ 1 }{ x }}(\frac{ x+1 }{ x ^{2} }-\frac{ 1 }{ x })\] wich is much easier to operate.

OpenStudy (bahrom7893):

btw the answer is -infinity

OpenStudy (owlcoffee):

I see, so it's -infinity and not 0 ?

OpenStudy (anonymous):

Since I made an error in the denominator, I would have to go back to the drawing board. But bahrom7893 says it is - infinity. Although I would like to through that again.